In Bohr’s series of lines of hydrogen spectrum, the third line from the red end corresponds to which of the following inter-orbit jumps of the electron for Bohr’s orbits in a hydrogen atom? why?
1)3 to 1
2)5 to 2
3)2 to 5
4)3 to 2

Grade:11

7 Answers

Vasantha Kumari
9 years ago

The correct answer is option b (5 to 2).

Here, we see red end means it is a part of the visible region and obviously only Balmer series corresponds to the visible region.

For Balmer series n1=2 and red end means low energy or third end from this end means n2=–5.

2®5

3

2®4

2

2®3

1

Transition

Line Number

Thanks & Regards,

Vasantha Sivaraj,

askIITians faculty

Jatin
6 years ago
Barmer series corresponds to the visible region.In barmer series n1=2 so it means the first line=3-2=1,second line means =4-2=2,third line=5-2=3
Jatin
6 years ago
the Lyman series it is the transitions from n >= 2 to the n = 1 state, for Balmer series its from n >= 3 to the n = 2 state and the Paschen series are transitions from n >= 4 to n = 3 state. The first line in each series is the transition from the next lowest number in the series to the lowest (so in the Lyman series the first line would be from n=2 to n=1) and the second line would be from from the third lowest to the lowest (in Lyman it would be n=3 to n=1) etc etc. To calculate the wavelength you can use the Rydberg formula.
kumar kanishk
6 years ago
The emission is in visible (VIBGYOR) region.Thus Balmer series start from n = 2 .The third line will be in the jump from 5 to 2.
Shumbul Arifa
5 years ago
Firstly , option 2-5 is eliminated since that requires energy but we want energy to be generated .Then , since they have clearly given that 3 lines are included , the option 3-2 is also eliminated. I hope you remember that only Balmer series is the one that involves emission in the visible spectrum .Thus , n2 must be 2 .. and since three lines are involved , n1 is 5 .Thank u
Deepak Dagar Jaat
5 years ago
Red end means in a question that to come visible region which is of balmer series 
n=2
third line means n=-5 so
ans. is 5 to 2
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Here, we see red end means it is a part of the visible region and obviously only Balmer series corresponds to the visible region.
For Balmer series n1=2 and red end means low energy or third end from this end means n2=–5.
2®5 3
2®4 2
2®3 1
Transition Line Number

Thanks and Regards